OpenStudy (anonymous):

Can somebody help me compare and contrast division of integers to division of rational expressions?

OpenStudy (anonymous):

ok so what is the question?

OpenStudy (anonymous):

That is the question...

OpenStudy (anonymous):

Compare and contrast division of integers to division of rational expressions.

OpenStudy (e.mccormick):

A rational expression is, basically, a polynomial over a polynomial. Like: \(\dfrac{x^2+9}{x-3}\)

OpenStudy (e.mccormick):

Division of integers is, or can be, a fraction: \(\dfrac{1}{2}\) or \(\dfrac{9}{2}\) and so on. That is what they want you to contrast.

OpenStudy (anonymous):

So could I say: “The difference between a rational expression and an integer is that in a rational expression you are dividing variables and polynomials while dividing integers you are just dividing whole numbers or fractions with no variables.”

OpenStudy (e.mccormick):

That would be a very reasonable way to say it.

OpenStudy (e.mccormick):

Hmmm, well, whole numbers are part of integers. So you might change the last part a bit.

OpenStudy (e.mccormick):

"...dividing integers you are just dividing whole numbers or fractions with no variables.” "...dividing integers you are just dividing the integers or making fractions with no variables.” See, that removes "whole numbers" which has a different meaning than integers.

OpenStudy (anonymous):

Sounds good, thanks :)